OpenStudy (anonymous):

What are the solutions of 6x2 = 12x? x = 0, x = −2 x = 0, x = 2 x = 6, x = 2 x = 6, x = −2

OpenStudy (anonymous):

6x^2 = 12x y = -6x^2 + 12x + 0 This is a quadratic equation. You can solve this with the quadratic formula:

OpenStudy (anonymous):

This is the formula.

OpenStudy (anonymous):

Just plug in the known values and solve. :)

OpenStudy (anonymous):

x = 0 x = 2

OpenStudy (anonymous):

-.- fine but idk what im plugging in!

OpenStudy (anonymous):

That's your answer. :)

OpenStudy (anonymous):

isnt it supposed to be -2?

OpenStudy (anonymous):

No, the answer is B.

jimthompson5910 (jim_thompson5910):

or... 6x^2 = 12x 6x^2 - 12x = 0 6x(x-2) = 0 6x = 0 or x-2 = 0 x = 0 or x = 2 which gives you the same answers

OpenStudy (anonymous):

alright, ohh i get it.. thank you both!

jimthompson5910 (jim_thompson5910):

the quadratic formula is useful in any case you have a quadratic...so it's probably best to get used to the formula (and use it anyway)

jimthompson5910 (jim_thompson5910):

so you can get practice

OpenStudy (anonymous):

You can check this by putting both of the answers in to the equation. 6(0)^2 = 12(0) 0 = 0 6(2)^2 = 12(2) 6(4) = 24 24 = 24 See? Both 0 and 2 are solutions.

OpenStudy (anonymous):

yes i do.. help on my other problem?